OpenStudy (anonymous):

Solve the equation. t – 32.8 = –27

OpenStudy (anonymous):

Add -32.8 to the right side of the equation and u have your answer

OpenStudy (anonymous):

add 32.8 not the negative

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

t-32.8=-27 t=-27+32.8

OpenStudy (anonymous):

add normally and you have your answer

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

t = -27 + 32.8 then t = ?

OpenStudy (anonymous):

do i get a medal?